Help2 Lavage Rapide is a Canadian company that owns and operates a large automatic car wash facility near Montreal. The following
Answer #1
Actual Revenue and Spending Variances Flexible budget
Car washed 8200 8200
Revenue 50700 (50700-49200)= 1500 F (8200*$6)= 49200
Expenses:
Cleaning supplies 4550 (4550-4100)= 450 U (8200*0.50)= 4100
Electricity 2199 (2199-2238)= 39 F (8200*0.09+1500)= 2238
Maintenance 1050 (1050-820)= 230 U (8200*0.10)= 820
Wages and salaries 6990 (6990-6660)= 330 U (8200*0.30+4200)= 6660
Depreciation 8200 0 NONE 8200
Rent 2200 (2200-2000)= 200 U 2000
Administration expenses 1762 (1762-1864)= 102 F (8200*0.02+1700)= 1864
Total expenses 26951 1069 U 25882
Net operating income $23749 431 F $23318
Lavage Rapide
Revenue and Spending Variances
For the Month Ended August 31
Revenue $1500 F
Expenses:
Cleaning supplies 450 U
Electricity 39 F
Maintenance 230 U
Wages and salaries 330 U
Depreciation 0 NONE
Rent 200 U
Administration expenses 102 F
Total expenses 1069 U
Net operating income $431 F